# Performance of a SiPM-based, plastic scintillator muon veto prototype for CUPID

### v1

@@ 3.1 SiPM quality assurance testing and coupling
Custom printed circuit boards (PCBs), shown in figure 4, were designed to hold 4 SiPMs per board for Prototypes 1 and 2. A ribbon cable supplies power to each SiPM and transmits its signal to the data acquisition (DAQ) system as described in section 3.2. The boards were installed inside cutouts in the plastic scintillator using custom 3D-printed plastic holders. A small layer of optical grease was applied to SiPMs to couple them to the WLS fiber ends.
The Small Panel SiPM [31] was soldered to an Adafruit Perma-Proto Small Tin board [32] and mounted to the Small Panel with a custom 3D-printed holder. The WLS fibers stuck out 1 cm from the panel, as shown in figure 3, and were held nearly flush against the SiPM by the mounted holder.
@@
Single photoelectron (p.e.) spectrum collected using the DAQ and LED driver.
The SiPMs were operated at 55 V (breakdown voltage Vbr 51 V) during all data collection, while the overvoltage was modified during QA to minimize noise and dark count rates. The light intensity of the LED driver was set to a value of 450 (arbitrary units) via the amplitude dial, and has a typical wavelength of 405 nm. The results of these tests are shown in figure 5 for a well-functioning SiPM. Currently, our setup suffers from the addition of noise due to inadequate grounding and noise coming from electronics, as it can be noticed in figure 5. A dedicated investigation to mitigate this noise is ongoing.
@@
A more rigorous QA procedure is being developed for use in the full muon veto system and has been performed on a sample of the SiPMs. In addition to the cable continuity and visual inspection tests described in the section above, this procedure includes measuring SiPM breakdown voltages after soldering and uses the CAEN V1730 digitizer [35] to obtain waveforms (our current readout does not contain a digitizer, c.f. section 3.2), Fourier transforms of those waveforms, and SPE spectra.
@@ 3.2 Electronics, readout, and DAQ
A CAEN DT5202 readout module was chosen for the muon veto data acquisition (DAQ) system [36], having single photoelectron energy resolution and timing resolution of 0.5 ns. This unit is designed to provide a bias voltage to SiPMs and read out the signals of a large array of detectors using a single cable. CAEN’s DT5202 readout software was modified to store the output in binary files for posterior use by the analysis framework Janus [37]. Information, such as integrated charge, trigger time, and timestamp, is stored in the output file. The CAEN DT5202 module contains two Citiroc-1A chips [38], each consisting of 32 channels, for a total of 64 channels. The relevant trigger modes are described below in brief; please consult the CAEN DT5202 manual for further detail [36].
The discriminator output for each the module’s 64 channels acts as a self-trigger, and the TLOGIC trigger source generates a trigger signal from logical combinations of these self-triggers. For Prototype 1, the 2.5 cm single panel prototype, the T-OR logic trigger is used. In this mode, an event is triggered when any of the active CAEN DT5202 channels self-trigger, and the output signals from all active channels are saved. The AND2_OR32 logic trigger is used for Prototype 2, which consists of a pair of 1 cm panels operating in coincidence. In this mode, an event is triggered whenever any pair of consecutive channels (e.g., channels 0 and 1, or channels 2 and 3, but not channels 1 and 2) from a Citiroc-1A chip have self-triggers that overlap by at least one clock cycle, or 8 ns.
@@
Self-trigger threshold values were determined from staircase scans over a range of threshold values performed with the Janus software. Staircase scans were completed with two different setups: SiPMs decoupled from the panel placed inside a dark box with an HV bias, and SiPMs coupled to the panels with no HV bias. From the first scan, a threshold value is chosen to minimize the event rate attributed to the intrinsic dark current at environmental temperature. The second scan helps to verify the DAQ is not triggering on electrical noise. The higher of the threshold values was set for all future runs taken with the panels.
@@ 3.3 Data taking
Data collection for these prototypes occurred at the Yale Wright Laboratory in New Haven, CT, which is 22.8 m above sea level. Panel prototypes were placed in custom dark boxes and covered with blackout fabric to reduce noise from ambient light in the SiPMs, see the resulting background spectra for each prototype in figure 7. SiPMs were connected to the DAQ via CAEN A5261 remotization cables [39] and AMPMODU type 3-102203-3 connectors [40].
@@ 4.1 Light yield
where ADCLG and ADCHG refer to the DAQ high gain and low gain readout, respectively (see section 3.2). The high gain and low gain are both needed for this calculation as the muon peak is visible only in the low gain (the relatively high energies of muon events get saturated in the high gain) and the high gain amplification is required to distinguish individual p.e. peaks.
The first term in equation 4.1 comes from the most probable value (MPV) from the Landau-Gaussian convolution fit to the muon peak in the low gain data (see figure 8b). The MPV is divided by the average energy deposited in a panel by a muon. This value was estimated by simulating our setup in Geant4 [41, 42] using a flux, obtaining a value of () MeV for a 2.5 cm thick panel. The second term is obtained by fitting the p.e. peaks in the high gain data with a series of Gaussians and averaging the difference in ADC between adjacent peaks. The fit was performed over a portion of the full ADC range that contained distinct peaks with similar amplitudes (see figure 8a). Variations in the difference between adjacent peaks arise from statistical fluctuations and are consistent across the full ADC range. Truncating the p.e. spectrum as shown in figure 8 is done for ease of fitting only, and the independence of fluctuations in the ADC difference between peaks was confirmed by initial fits of p.e. peaks over both the full range and various other truncated portions of the high gain data. Finally, the third term represents the ratio between the high gain and low gain ADC from calibration data. This comes from plotting the high gain data against the low gain data and performing a linear fit. Only events that do not saturate the DAQ’s high gain mode are used for this calculation, and the linear trend is extrapolated to higher energy events (see figure 8).
@@
To account for non-homogeneity between mini-modules in Prototype 1, the quantities in equation 4.1 are calculated for each mini-module individually. The values of each readout channel are summed and the two latter terms are determined by a weighted average based on each mini-module’s contribution to the total . The homogeneity of the mini-modules themselves was confirmed by operating the Small Panel in coincidence with a 2562.5 cm3 piece of plastic scintillator with a single WLS fiber and SiPM. The performance of the Small Panel was consistent for muon events in each region, indicating that the mini-module design provides sufficient homogeneity. We found light yields of () p.e./MeV for the Small Panel and () p.e./MeV for the full Prototype 1 panel. The light yield was not calculated for the Prototype 2 panels because the 1 cm thickness does not yield a separation between the gamma and muon peaks in the ADCLG spectrum that is sufficient for the calculation to be reliable.
4.2 Light yield uniformity
@@
To characterize the light yield uniformity of the LBNL panel, Prototype 1, and Prototype 2, a collimated 60Co source was placed at the center of each mini-module of the Prototype 1 and 2 panels, and at 16 uniformly-spaced positions across the LBNL panel. The counts were normalized to the highest number of counts for the module being tested, so the comparison between different prototypes is fair. For Prototype 2, a single 1 cm thick panel was characterized. Data was taken for 5 minutes at each position of the source for the LBNL panel and for 2 minutes for Prototypes 1 and 2. The variations in light yield for each location of the 60Co source are shown in figure 9.
@@
The summation represents the total number of 60Co events observed by all mini-modules the 2-minute data-taking interval. The LBNL panel is not physically segmented, so the summation represents the number of events detected by the PMT for the whole panel. Although this summation includes muon and gamma events, we can assume their respective fluxes to be uniform over time and to not represent a source of bias in our analysis. The source location with the highest summed ADC counts was used to normalize all other locations. The statistical error on all of the data points is less than 1%. For Prototype 1, the light yield was the most uniform. The average normalized number of 60Co events was 0.97 0.02, where the error is the statistical uncertainty. For Prototype 2, the uniformity is slightly worse with an average of 0.87 0.06. The least uniform light yield was observed in the LBNL panel with an average of 0.83 0.09.
@@
The behavior of light yield changes in a structured way for the LBNL panel due to the uniformity of the design. The light yield is highest for locations closest to the PMT and the light yield worsens for locations along the edge of the panel. The motivation for the swirly design of the fibers in Prototypes 1 and 2 was to increase light yield uniformity which is what we observe. This allows these panels to have higher efficiency in tagging muons. However, this introduces a source of systematics that affects our panels uniquely— the alignment of the wavelength-shifting fiber with the SiPMs. While we try to minimize this effect by the use of a stencil during fabrication, the small area of the SiPM device makes it challenging. Since the diameter of each of the two fibers are 1 mm, and they feed into a 33 mm2 SiPM, there is little room for error in their alignment. This variation can contribute to the light yield efficiency within a panel as well as across panels.
@@ 4.3 Muon discrimination power
Muon discrimination refers to the ability to differentiate between muon and gamma events in the detector, as these distributions overlap in the energy spectrum for our system. For a given (low gain) ADC threshold, events with a greater ADC value (right of the threshold) are tagged as muons. We estimate the fraction of all muons that get tagged using a data-driven approach in which the gamma peaks follow a Gaussian distribution whereas the muon events are distributed according to a Landau-Gaussian distribution, as shown in figure 10.
The muon tagging efficiency is obtained by integrating the Landau-Gaussian distribution from a chosen ADC threshold, which separates the two distributions, and normalizing by the marginalized Landau-Gaussian distribution. We chose this threshold to be 1600 ADC. We find that the ratio of muons that are tagged above that threshold is . Similarly, for that particular threshold choice, the false positive rate from gamma events corresponds to 0.34 Hz for a self-triggered single module. To estimate the induced deadtime, we assume a similar false-positive rate underground at LNGS and an estimated covered area of 78 m2 by the complete muon veto system. We expect an induced deadtime of in CUPID for a veto window of 0.5 ms (CUPID’s light detector rise-time [4]) by assuming coincidence hits between two panels intersected by the muon track. This meets CUPID’s deadtime requirement of , and it is mainly dominated by false-positive events from the gamma peak leaking into the muon peak, making the induced rate from real muon events negligible.
@@ 4.4 Detection efficiency
We estimate the detection efficiency, i.e. the fraction of muons detected (above threshold) compared to the total number of muons that pass through the single module. To do this, we "sandwich" the Prototype 1 module between two panels that trigger in coincidence, while the channels in the single module are masked out and do not participate in the triggering scheme. The panels are physically separated as shown in figure 11, for which the separation was arbitrarily chosen and determined mainly by the dimensional constraints of our dark-box.
With this configuration, we estimate the detection efficiency to be . We assume a detection efficiency of for the coincidence panels, so we do not apply any corrections to the final number we report here. Although this efficiency is high and consistent with other similar designs [16], it can be further improved by increasing the thickness of the scintillator panel or operating it as a coincidence using two stacked modules instead of one. This might be necessary for ultra-low background experiments with stringent background budgets.
@@ 4.5 Position reconstruction of muon events
One motivation for including eight readouts in each panel of Prototypes 1 and 2 was to enable coarse position reconstruction of events. To quantify the position reconstruction of Prototype 1 for muon events, eight 30-minute runs were taken using the Small Panel prototype (2525 cm2 panel) as a coincidence panel with Prototype 1 in various locations following the same location labeling as figure 9. OR trigger logic was used for data taking, and offline coincidence events were found between the Small Panel and Prototype 1, as shown in figure 3. In this analysis, muon events were considered to be coincident events above a defined threshold in the Small Panel’s energy spectra above the gamma region. Because the Small Panel rests on top of Prototype 1, the position of a muon event was assumed to be in the same area as the location of the Small Panel for each data run.
@@
For each location, the spectrum from the eight SiPMs, the Small Panel, and the sum of all SiPM channels converted to p.e. were examined. In most locations, the SiPM channel under the location of the Small Panel has the highest p.e. value compared to other channels. This suggests that the sub-section of the panel the muon passes through often has the highest light yield. With these observations, two simple algorithms were designed to identify the location of muon events and estimate position reconstruction capabilities. The first algorithm assigned the locations of muon events as the mini-modules containing the SiPM channel with the highest p.e. count in an event, which covers a 2525 cm2 area. The algorithm was also used to save the highest two SiPM channels to reconstruct the event to a 2550 cm2 area. These algorithms were run on the eight collected datasets for the single module (before any coincidence cuts) to test their efficiencies since the muon events in each location can be isolated and are known. The fraction of muon events that these algorithms correctly reconstructed the position of is shown in figure 12. The track reconstruction algorithm using the highest two channels had an efficiency of and the single track channel reconstruction was .
@@
Additional position reconstruction algorithms could be developed for track reconstruction using more SiPM channels and metrics within a panel such as the ratio of ADC counts in an event, multiple panels, and information from events in CUORE crystals. This development is beyond the scope of this work. However, from the simple algorithms described, we estimate that using Prototype 1, the position of muon events is expected to reconstruct to 2525 cm2 with approximately an 80 efficiency. Because Prototype 2 operates as a double panel module between mini-modules of 2525 cm2, its position reconstruction is assumed to be 2525 cm2 from the two adjacent channels that trigger in coincidence mode on the top and bottom panels. The possibility for finer position reconstruction is an advantage of including multiple mini-modules with SiPM light detectors in one larger module instead of using one light detector per module.

### v2

@@ 3.1 SiPM coupling and quality assurance
Custom printed circuit boards (PCBs), shown in figure 4, were designed to hold 4 SiPMs each for Prototypes 1 and 2. A ribbon cable supplies power to each SiPM and transmits its signal to the data acquisition (DAQ) system as described in section 3.2. The boards are installed inside cutouts in the plastic scintillator using custom 3D-printed plastic holders. A small layer of Eljen Technologies EJ-550 silicon optical grease wass applied to the SiPMs to couple them to the ends of the WLS fibers [35].
The Small Panel uses one 13360-3050 Hamamatsu SiPM soldered to an Adafruit Perma-Proto Small Tin board [36] and mounted to the Small Panel with a custom 3D-printed holder. The WLS fibers protrude 1 cm from the panel, as shown in figure 3, and are held nearly flush against the SiPM by the mounted holder.
@@
Single photoelectron spectrum collected using the DAQ and LED driver.
The SiPMs were operated at 55 V (breakdown voltage Vbr 51 V) during all data collection, while the overvoltage was modified during QA to minimize noise and dark count rates. The light intensity of the LED driver was set to a value of 450 (arbitrary units) via the amplitude dial, and has a typical wavelength of 405 nm. The results of these tests are shown in figure 5 for a well-functioning SiPM. Currently, our setup suffers from the addition of noise due to inadequate grounding and electronic interference, as can be seen in figure 5. A dedicated investigation to mitigate this noise is ongoing.
@@
A more rigorous QA procedure is being developed for use in the full muon veto system and has been performed on a sample of the SiPMs. In addition to the cable continuity and visual inspection tests described in the section above, this procedure includes measuring SiPM breakdown voltages after soldering and uses the CAEN V1730 digitizer [38] to obtain waveforms (our current readout does not contain a digitizer, cf. section 3.2), Fourier transforms of those waveforms, and single photoelectron spectra.
@@ 3.2 Electronics, readout, and DAQ
A CAEN DT5202 readout module was chosen for the muon veto data acquisition (DAQ) system [40], having single photoelectron energy resolution and a timing resolution of 0.5 ns. This unit is designed to provide a bias voltage to SiPMs and read out the signals of a large array of detectors using a single cable. CAEN’s DT5202 readout software was modified to store the output in binary files for posterior use by the analysis framework Janus [41]. Information such as integrated charge, trigger time, and timestamp is stored in the output file. The CAEN DT5202 module contains two Citiroc-1A chips [42], each consisting of 32 channels, for a total of 64 channels. The relevant trigger modes are described below in brief; please consult the CAEN DT5202 manual for further detail [40].
The discriminator output for each the module’s 64 channels acts as a self-trigger, and the TLOGIC trigger source generates a trigger signal from logical combinations of these self-triggers. For Prototype 1, the 2.5 cm single panel module, the T-OR logic trigger is used. In this mode, an event is triggered when any of the active CAEN DT5202 channels self-trigger, and the output signals from all active channels are saved. The AND2_OR32 logic trigger is used for Prototype 2, which consists of a pair of 1 cm panels operating in coincidence. In this mode, an event is triggered whenever any pair of consecutive (even, odd)-ordered channels (e.g., channels 0 and 1, or channels 2 and 3, but not channels 1 and 2) from a Citiroc-1A chip have self-triggers that overlap by at least one clock cycle, or 8 ns. The panel on the top in Prototype 2 is connected even channels (0, 2, …, 14) and the bottom panel is connected to odd channels (1, 3, …, 15).
@@
Self-trigger threshold values were determined from staircase scans over a range of threshold values performed with the Janus software. Staircase scans were completed with two different setups: SiPMs decoupled from the panel placed inside a dark box with a reverse bias voltage, and SiPMs coupled to the panels with no bias voltage. From the first scan, a threshold value is chosen to minimize the event rate attributed to the intrinsic dark current at environmental temperature. The second scan helps to verify the DAQ is not triggering on electrical noise. The higher of the threshold values was set for all future runs taken with the panels.
@@ 3.3 Data taking
Data collection for these prototypes occurred at the Yale Wright Laboratory in New Haven, CT, which is 22.8 m above sea level. Panel prototypes were placed in custom dark boxes and covered with blackout fabric to reduce noise from ambient light in the SiPMs, see the resulting background spectra for each prototype in figure 7. This light tightness will be a requirement for the official installation at LNGS as artificial illumination will be present onsite. SiPMs were connected to the DAQ via CAEN A5261 remotization cables [43] and AMPMODU type 3-102203-3 connectors [44].
@@ 4.1 Light yield
where ADCLG and ADCHG refer to the DAQ low gain and high gain readout, respectively (see section 3.2). The high gain and low gain are both needed for this calculation as the muon peak is visible only in the low gain (the relatively high energies of muon events get saturated in the high gain) and the high gain amplification is required to distinguish individual p.e. peaks.
The first term in equation 4.1 comes from the most probable value (MPV) from the Landau-Gaussian convolution fit to the muon peak in the low gain data (see bottom left of figure 8). The MPV is divided by the average energy deposited in a panel by a muon. This value was estimated by simulating our setup in Geant4 [45, 46] using a flux distribution, which yielded a value of () MeV for a 2.5 cm thick panel.
The second term is obtained by fitting the p.e. peaks in the high gain data with a series of independent Gaussians and averaging the difference in ADC between adjacent peaks. The fit is performed over a portion of the full ADCHG range that contains distinct peaks with relatively similar amplitudes (see top left of figure 8). Variations in the difference between adjacent peaks arise from statistical fluctuations and are consistent across the full ADCHG range. The independence of these fluctuations was confirmed by performing fits of the full range of visible p.e. peaks and various sub-ranges. The fit shown in figure 8, for example, includes five p.e. peaks over the range [750, 1400] ADCHG and yields a value of ADCHG/p.e. Extending the fit to include the first 11 p.e. peaks over the range [470, 1990] ADCHG (this does not include the first peak at 0 ADCHG which corresponds to 0 p.e. (pedestal) events) yields a value of ADCHG/p.e., which is consistent within uncertainty.
Finally, the third term represents the ratio between the high gain and low gain ADC from calibration data. This comes from plotting the high gain data against the low gain data and performing a linear fit. Only events that do not saturate the DAQ’s high gain mode are used for this calculation, and the linear trend is extrapolated to higher energy events (see figure 8).
@@
To account for non-homogeneity between mini-modules in Prototype 1, the quantities in equation 4.1 are calculated for each mini-module individually. The values of each readout channel are summed and the two latter terms are determined by a weighted average based on each mini-module’s contribution to the total of the panel. The homogeneity of the mini-modules themselves was confirmed by operating the Small Panel in coincidence with a 2562.5 cm3 piece of plastic scintillator with a single WLS fiber and SiPM. The performance of the Small Panel was consistent for muon events in each region, indicating that the mini-module design provides sufficient homogeneity. We found a light yield of () p.e./MeV for the Small Panel and of () p.e./MeV for the full Prototype 1 panel. The light yield was not calculated for the Prototype 2 panels because the 1 cm thickness does not produce a separation between the gamma and muon peaks in the ADCLG spectrum that is sufficient for the calculation to be reliable.
4.2 Light collection uniformity
@@
To characterize the light collection uniformity of the LBNL Panel, Prototype 1, and Prototype 2, a collimated 60Co source is placed at the center of each mini-module of the Prototype 1 and 2 panels, and at 16 uniformly-spaced positions across the LBNL Panel. For Prototype 2, a single 1 cm thick panel was characterized. Data was taken for 5 minutes at each position of the source for the LBNL Panel and for 2 minutes for Prototypes 1 and 2. The ADCLG output from each source location is normalized to that of the source location that produced the highest number of counts in a given module, allowing for comparison between prototypes. The variations in light collection for each location of the 60Co source are shown in figure 9.
@@
To quantify the uniformity of the different panels, the total ADCLG spectrum is integrated for each data-taking interval (i.e., for each source location the total output from the panel was summed). The summation represents the total number of events observed by all mini-modules in the data-taking interval. The LBNL Panel is not physically segmented, so the summation represents the number of events detected by the PMT for the whole panel. Although this summation includes muon and gamma events in addition to the 60Co events, we can assume their respective fluxes to be uniform over time and to not represent a source of bias in our analysis. As described above, the source location with the largest sum of ADCLG counts is used to normalize all other locations. The statistical error is less than 1% for all data points. Prototype 1 has the most uniform light collection. The average normalized number of 60Co events is 0.97 0.02, where the error is the statistical uncertainty. For Prototype 2, the uniformity is slightly worse with an average of 0.87 0.06 normalized 60Co events. The least uniform light collection was observed in the LBNL Panel with an average of 0.83 0.09 normalized 60Co events.
@@
The behavior of light collection changes in a structured way for the LBNL Panel due to the unsegmented design. The light collection is highest for locations closest to the PMT and the light collection worsens for locations along the edge of the panel. The motivation for the swirly design of the fibers in Prototypes 1 and 2 was to increase the light collection uniformity, which is what we observe. This allows these panels to have higher efficiency in tagging muons. However, this introduces a source of systematics that affects our panels uniquely: the alignment of the WLS fibers with the SiPMs. Although we tried to minimize this effect by using a stencil during fabrication, the small area of the SiPM made this challenging. Since the diameter of each of the two fibers are 1 mm, and they feed into a 33 mm2 SiPM, there is little room for error in their alignment. This variation can contribute to the light collection fluctuation within a panel, as well as across panels.
@@ 4.3 Muon discrimination power
Muon discrimination refers to the ability to differentiate between muon and gamma events, whose distributions overlap in our panels’ energy spectra. For a given (low gain) ADC threshold, events with a greater ADC value (to the right of the threshold) are tagged as muons. We estimate the fraction of all muons that get tagged using a data-driven approach in which the gamma peaks follow a Gaussian distribution whereas the muon events are distributed according to a Landau-Gaussian distribution, as shown in figure 10.
The muon tagging efficiency is obtained by integrating the Landau-Gaussian distribution from a chosen ADC threshold, which separates the two distributions, and normalizing by the marginalized Landau-Gaussian distribution. We chose this threshold to be 1600 ADC. We find that the ratio of muons that are tagged above that threshold for Prototype 1 is . Similarly, for that particular threshold choice, the false positive rate from gamma events corresponds to 0.34 Hz for a self-triggered single module. To estimate the induced deadtime, we conservatively [47, 48] assume a similar false-positive rate underground at LNGS, and an estimated coverage area of 78 m2 by the complete muon veto system. We expect an induced deadtime of in CUPID for a veto window of 0.5 ms (CUPID’s light detector rise-time [4]) by assuming coincidence hits between two panels intersected by the muon track. This meets CUPID’s deadtime requirement of , and is mainly dominated by false-positive events from the gamma peak leaking into the muon peak, making the induced rate from real muon events negligible. This estimation was not performed for Prototype 2 since the gamma/muon separation is not resolvable.
@@ 4.4 Detection efficiency
We estimate the detection efficiency, i.e. the fraction of muons detected (above a chosen threshold, see Fig. 10) compared to the total number of muons that pass through the single module. To do this, we "sandwich" the Prototype 1 module between the two Prototype 2 panels, which trigger in coincidence, while the channels in the Prototype 1 module are masked out and do not participate in the triggering scheme. The panels are physically separated as shown in figure 11, with an arbitrarily chosen separation length determined mainly by the dimensional constraints of our dark box.
With this configuration, we estimate the detection efficiency of Prototype 1 to be . We assume a detection efficiency of for Prototype 2, so we do not apply any corrections to the final number that we report here. Although this efficiency is already high and consistent with other similar designs [18], it can be further improved by increasing the thickness of the scintillator panel or operating in coincidence using two stacked modules instead of one. This might be necessary for ultra-low background experiments with stringent background budgets.
@@ 4.5 Position reconstruction of muon events
One motivation for including eight readouts in each panel of Prototypes 1 and 2 was to enable coarse position reconstruction of events. To quantify the position reconstruction of Prototype 1 for muon events, eight 30-minute runs were taken using the Small Panel prototype (2525 cm2 panel) in a coincidence configuration with the Prototype 1 mini-modules, following the same location labeling as figure 9. OR trigger logic was used for data taking, and coincidence events were found offline between the Small Panel and Prototype 1. In this analysis, muon events were considered to be coincidence events above a defined (low gain) ADC threshold in the Small Panel’s energy spectrum. Because the Small Panel rests directly on top of Prototype 1, the position of a muon event is assumed to be in the same area as the location of the Small Panel for each data run.
@@
For each location, the spectrum (in p.e.) from each of the Prototype 1 SiPM channels, the Small Panel, and the sum of Prototype 1 channels were examined. In most locations, the SiPM channel with the highest muon peak p.e. in Prototype 1 corresponds to the mini-module directly under the Small Panel. This suggests that the sub-section of the panel the muon passes through typically has the highest light yield. With these observations, two simple algorithms were designed to identify the location of muon events and estimate position reconstruction capabilities. The first algorithm assigns the location of a muon event to the mini-module channel with the highest muon peak p.e., which covers a 2525 cm2 area. The second algorithm assigns both the mini-module channel with the highest and the channel with the second highest muon p.e. peak to each muon event, reconstructing the event to a 2550 cm2 area. These algorithms were run on the eight collected datasets for the Prototype 1 module (before any coincidence cuts) to test their efficiencies as the muon events in each location can be isolated and are known. The fraction of correctly reconstructed muon events is shown in figure 12. The track reconstruction algorithm using the highest two channels had an efficiency of , and the single track channel reconstruction had an efficiency of .
@@
Additional position reconstruction algorithms could be developed for track reconstruction using more SiPM channels and metrics within a panel such as the ratio of ADC counts in an event, multiple panels, and information from events in CUORE crystals. This development is beyond the scope of this work. However, from the simple algorithms described, we estimate that using Prototype 1, the position of muon events is expected to reconstruct to 2525 cm2 with approximately an 80 efficiency. Because Prototype 2 operates as a double panel module between mini-modules of 2525 cm2, its position reconstruction granularity is assumed to be 2525 cm2 from the two adjacent channels that trigger in coincidence on the top and bottom panels. The possibility for finer position reconstruction is an advantage of including multiple mini-modules with their own light detectors in one larger module instead of using one light detector per module.